# Self-Consistent Stochastic Finite-Temperature Modelling: Ultracold Bose Gases with Local (s-wave) and Long-Range (Dipolar) Interactions

### v1

@@ V.4.1 Consistency of the LHY Correction Term
Before closing this stochastic section we add some further remarks that may have some potential relevance here: A related detailed comparison between different formalisms, most notably between kinetic-type (ZNG) and stochastic-based (SGPE) approaches, has also been elusive in the case of atomic gases with contact interactions. [^12] See, e.g., Chapters discussing such connections by Griffin and Zaremba, and Wright, Davis and Proukakis in Ref. [18]. Some interesting advances were nonetheless made through tailored numerical simulations in the framework of the Projected GPE [146, 65] and stochastic GPE settings: specifically, in [196, 86] finite-temperature equilibrium data obtained respectively within the PGPE and SGPE were used to extract the condensate contribution through the Penrose-Onsager mode criterion [197] (corresponding to the mode with largest eigenvalue). Number-conserving normal and anomalous averages of fluctuation operators beyond the numerically-obtained Penrose-Onsager mode were the constructed, and their relative importance ascertained (e.g. by studying their effect on the ‘renormalized’ equilibrium system chemical potential, which should remain independent of position in a homogeneous system [196]). A key takeaway from such works is that when using simulations with noisy fields (including, e.g., noisy initial conditions in PGPE, dynamical noise in SGPE), anomalous averages of carefully-constructed beyond-condensate number-conserving operators – which are likely to contain many-body effects, and quantum depletion features related to the LHY correction – can in fact, at least to some extent, arise naturally within such formalisms. In other words, the interpretation of what constitutes the ‘condensate’ within such multi-mode PGPE/SGPE treatments (i.e. contrasting the full $|\Phi_{0}|^{2}$ field to the self-consistently determined Penrose-Onsager mode) could imply that the appearance of the LHY correction in the dynamical generalised GPE equations should not necessarily be in the usual form of $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$, as $\Phi_{0}$ is explicitly a multi-mode field spanning not only the condensate itself, but also the low-lying modes affected by its presence. Instead, it is plausible that a term with the usual LHY structure should only be written in terms of the corresponding numerically-obtained Penrose-Onsager mode, $\Phi_{PO}$. If such an interpretation were correct, this could somewhat question the premise of using the dipolar SPGPE in the above-stated form of Eq. (87) (without any further amendments). Interestingly, such an issue need not however necessarily arise in the ZNG-like treatment of Sec. V.3, since there the field $\Phi_{0}$ refers explicitly to the condensate mode. While such considerations may in fact have very little effect in practice in the numerical calculations and thus any subsequent comparison to experiments, we nonetheless felt it was pertinent to comment on this here, at least from a fundamental perspective, for completeness.
VI Conclusions
@@
In this work we have presented a general non-equilibrium theoretical formalism for modelling Bose gases with a general interatomic potential, which is expected to be valid in a broad range of physical scenarios: while we have explicitly demonstrated their relevance in bridging existing theories and extending beyond them in the context of ultracold Bose gases with contact interactions, and dipolar gas condensates, we anticipate that the more general formulation of our theory (discussed in Appendix A) might in the future also become useful in modelling other systems exhibiting long-range interactions, such as dipolar molecular condensates, or Rydberg systems. Our formalism is based on splitting the system into coherent and incoherent degrees of freedom, and including fluctuations of the incoherent modes to all orders, but fluctuations in the coherent modes are only maintained to second order in the Keldysh q-fields, representing up to stochastic (statistical) fluctuations. As such, our approach is valid in the domain when thermal fluctuations dominate the system properties, where it was shown to generalize a range of established theories, such as the Gross-Pitaevskii-Boltzmann (or ZNG) model, and the stochastic (and projected) Gross-Pitaevskii model – both of which have found immense success in modelling ultracold atomic experiments with contact interactions.
@@
Motivated by recent discoveries in the context of dipolar gases, where regimes exist in which quantum fluctuations become so dominant, that they help stabilize the system against collapse acting as an effective repulsive force, we have given a brief review of how such distinct regime can be handled theoretically. Having an accurate description of the two limiting cases, namely quantum-dominated and thermal-dominated, has enabled us – following similar lines as a handful of recent works [192, 191] – to postulate a more generalised model which includes both quantum and thermal fluctuations of the coherent field, and may thus reasonably interpolate between the established limiting cases. In so doing, we have also formulated a potentially-relevant criterion [Eq. (76)] for the relative importance of quantum and thermal fluctuations in a given system. While the extent of such an extrapolation to the regime where quantum and thermal fluctuations co-exist in comparable measures is unknown, intuition from similar questions in ultracold gases with contact interactions and preliminary evidence from successfully modelling of dipolar gas experiments, provide promising avenues for further exploration with such combined approaches.
@@
Under such an assumption, our final set of equations for non-equilibrium finite-temperature systems exhibiting long-range interactions takes the form of the coupled Eqs. (12)-(16), but with Eq. (12) featuring an additional term accounting for the Lee-Huang-Yang quantum correction term. In the case of dipolar gases, such a term would take the usual form of Eq. (68), possibly even through its finite-temperature generalized expression also involving thermal densities, which in our treatment can be fully dynamically included through the quantum Boltzmann equation (15) for the phase-space distribution of the incoherent particles.
We hope this work will pave the way forward for more systematic studies of non-equilibrium finite-temperature quantum gases with long-range interactions, and look forward to further developments helping settle some of the questions raised within this work both on the theoretical side, and in terms of their applicability to regimes of fundamental and experimental relevance.
@@ Acknowledgements
This work was supported by the Leverhulme Trust, Grant no. RPG-2021-010. We acknowledge insightful discussions with Tom Bland, Ashton Bradley and Mike Garrett.
@@ Appendix A Final Stochastic Equations for a General Interatomic Potential
    $\alpha_{1}=g-\frac{C_{dd}}{3},\quad\quad\quad\alpha_{2}=C_{dd},\quad\quad\quad{\rm and}\quad\quad\quad\mathcal{O}=C_{dd}\,\frac{\nabla^{2}}{(\mathbf{n}\cdot\nabla)^{2}}\;.$ (106)
@@
    $\alpha_{1}=g,\quad\quad\quad\quad\quad\alpha_{2}=0\;.$ (107)
@@
    $\alpha_{1}=g,\quad\quad\quad\quad\quad\alpha_{2}=m,\quad\quad\quad\quad{\rm and}\quad\quad\quad\quad\mathcal{O}=\frac{1}{4\pi G}\,\nabla^{2}\;.$ (108)

### v2

@@ VI.4.1 Consistency of the LHY Correction Term
Before closing this stochastic section we add some further remarks that may have some potential relevance here: A related detailed comparison between different formalisms, most notably between kinetic-type (ZNG) and stochastic-based (SGPE) approaches, has also been elusive in the case of atomic gases with contact interactions [^19] See, e.g., Chapters by Griffin and Zaremba on the one hand, and by Wright, Davis and Proukakis on the other discussing such connections in Ref. [78].. Some interesting advances were nonetheless made through tailored numerical simulations in the framework of the Projected GPE [205, 124] and stochastic GPE settings: specifically, in [250, 145] finite-temperature equilibrium data obtained respectively within the PGPE and SGPE were used to extract the condensate contribution through the Penrose-Onsager mode criterion [251] (corresponding to the mode with largest eigenvalue). Number-conserving normal and anomalous averages of fluctuation operators beyond the numerically-obtained Penrose-Onsager mode were the constructed, and their relative importance ascertained (e.g. by studying their effect on the ‘renormalized’ equilibrium system chemical potential, which should remain independent of position in a homogeneous system [250]). A key takeaway from such works is that when using simulations with noisy fields (including, e.g., noisy initial conditions in PGPE, dynamical noise in SGPE), anomalous averages of carefully-constructed beyond-condensate number-conserving operators – which are likely to contain many-body effects, and quantum depletion features related to the LHY correction – can in fact, at least to some extent, arise naturally within such formalisms. In other words, the interpretation of what constitutes the ‘condensate’ within such multi-mode PGPE/SGPE treatments (i.e. contrasting the full $|\Phi_{0}|^{2}$ field to the self-consistently determined Penrose-Onsager mode) could imply that the appearance of the LHY correction in the dynamical generalised GPE equations should not necessarily be in the usual form of $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$, as $\Phi_{0}$ is explicitly a multi-mode field spanning not only the condensate itself, but also the low-lying modes affected by its presence. Instead, it is plausible that a term with the usual LHY structure should only be written in terms of the corresponding numerically-obtained Penrose-Onsager mode, $\Phi_{PO}$. If such an interpretation were correct, this could somewhat question the premise of using the dipolar SPGPE in the above-stated form of Eq. (88) (without any further amendments). Interestingly, such an issue need not however necessarily arise in the ZNG-like treatment of Sec. VI.3, since there the field $\Phi_{0}$ refers explicitly to the condensate mode. While such considerations may in fact have very little effect in practice in the numerical calculations and thus any subsequent comparison to experiments, we nonetheless felt it was pertinent to comment on this here, at least from a fundamental perspective, for completeness.
VII Conclusions
@@
In this work we have presented a general non-equilibrium theoretical formalism for modelling Bose gases with a general interatomic potential, which is expected to be valid in a broad range of physical scenarios: while we have explicitly demonstrated their relevance in bridging existing theories and extending beyond them in the context of ultracold Bose gases with contact interactions, and dipolar gas condensates, we anticipate that the more general formulation of our theory (discussed in Appendix A) might in the future also become useful in modelling other systems exhibiting long-range interactions, such as dipolar molecular condensates, Rydberg systems, or indirect excitons. Our formalism is based on splitting the system into coherent and incoherent degrees of freedom, and including fluctuations of the incoherent modes to all orders, but fluctuations in the coherent modes are only maintained up to second order in the Keldysh q-fields, representing stochastic (statistical) fluctuations. As such, our approach is valid in the domain when thermal fluctuations dominate the system properties; in such a limit, our formalism was shown to generalize a range of established theories, such as the Gross-Pitaevskii-Boltzmann (or ZNG) model, and the stochastic (and projected) Gross-Pitaevskii model – both of which have found immense success in modelling ultracold atomic experiments with contact interactions.
@@
Motivated by recent discoveries in the context of dipolar gases, where regimes exist in which quantum fluctuations become dominant, helping stabilize the system against collapse, we have given a brief review of how such distinct regime could potentially be handled theoretically. Having an accurate description of the two limiting cases, namely quantum-dominated and thermal-dominated, has enabled us – following similar lines as a handful of recent works [57, 47] – to postulate a more generalised model which includes both quantum and thermal fluctuations of the coherent field, and may thus reasonably interpolate between the established limiting cases. In so doing, we have also formulated a potentially-relevant criterion [Eq. (77)] for the relative importance of quantum and thermal fluctuations in a given system. While the extent of such an extrapolation to the regime where quantum and thermal fluctuations co-exist in comparable measures is unknown, intuition from similar questions in ultracold gases with contact interactions and preliminary evidence from successfully modelling of dipolar gas experiments, provide promising avenues for further exploration with such combined approaches.
@@
Under such an assumption, our final set of equations for non-equilibrium finite-temperature systems exhibiting long-range interactions takes the form of the coupled Eqs. (12)-(17), with the explicit inclusion of an additional term accounting for the Lee-Huang-Yang quantum correction. In the case of dipolar gases, such a term would take the usual form of Eq. (69), possibly even through its finite-temperature generalized expression also involving thermal densities, which in our treatment can be fully dynamically included through the quantum Boltzmann equation (16) for the phase-space distribution of the incoherent particles.
While admittedly various open issues remain, we believe this work represents another important step towards the development of systematic studies of non-equilibrium finite-temperature quantum gases with long-range interactions.
@@ Acknowledgements
This work was supported by the Leverhulme Trust, Grant no. RPG-2021-010. We acknowledge insightful discussions with Tom Bland, Ashton Bradley and Mike Garrett, and useful comments by Axel Pelster and Duncan O’Dell.
@@ Appendix A Final Stochastic Equations for a General Interatomic Potential
    $\hskip-56.9055pt\alpha_{1}=g-\frac{C_{dd}}{3},\quad\quad\quad\alpha_{2}=C_{dd},\quad\quad\quad{\rm and}\quad\quad\quad\mathcal{O}=C_{dd}\,\frac{\nabla^{2}}{(\mathbf{n}\cdot\nabla)^{2}}\;.$ (107)
@@
    $\hskip-221.93158pt\alpha_{1}=g,\quad\quad\quad\quad\quad\alpha_{2}=0\;.$ (108)
@@
    $\hskip-42.67912pt\alpha_{1}=g,\quad\quad\quad\quad\quad\alpha_{2}=m,\quad\quad\quad\quad{\rm and}\quad\quad\quad\quad\mathcal{O}=\frac{1}{4\pi G}\,\nabla^{2}\;.$ (109)